OpenStudy (calculusxy):

Clementine is 5 years younger than Matilda. In 15 years, their combined ages will be 45. How old is Clementine now?

jimthompson5910 (jim_thompson5910):

C = Clementine age now M = Matilda's age now

jimthompson5910 (jim_thompson5910):

"Clementine is 5 years younger than Matilda" So C = M-5

jimthompson5910 (jim_thompson5910):

15 yrs from now, the two ages will be C+15 and M+15 "In 15 years, their combined ages will be 45" which means (M+15) + (C+15) = 45 (M+15) + (M-5+15) = 45 ... plug in C = M-5 and solve for M

OpenStudy (calculusxy):

How do I solve for M?

jimthompson5910 (jim_thompson5910):

(M+15) + (M-5+15) = 45 M+15 + M-5+15 = 45 2M+25 = 45 I'll let you finish

OpenStudy (calculusxy):

So M=10?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (calculusxy):

So now do I do 10-5=5 (Clementine)

jimthompson5910 (jim_thompson5910):

yes C = 5